# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ 3.2 Implicit bias of mirror flow with scaled potentials
Assume the potential function $\Phi\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}$ satisfies: (i) $\Phi$ can be written in the form of $\Phi(\theta)=\frac{1}{n^{2}}\sum_{k=1}^{3n+1}\phi\big{(}n(\theta_{k}-\hat{\theta}_{k})\big{)}$, where $\phi$ is a real-valued function on $\mathbb{R}$; (ii) $\phi$ takes the form of $\phi(x)=\frac{1}{1+\omega}\big{(}\psi(x)+\omega x^{2}\big{)}$, where $\omega>0$ and $\psi\in C^{3}(\mathbb{R})$ is a convex function on $\mathbb{R}$.
@@
The following result characterizes the implicit bias of mirror with scaled potentials for wide networks with absolute value activations. We consider absolute value activations because, compared to ReLU they yield a smaller function space that makes the problem more tractable.
@@
Consider a two layer network (1) with absolute value activation and $n$ hidden units. Assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Additionally, assume $\mathcal{B}$ is compactly supported. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with a potential satisfying Assumptions 7 and learning rate $\eta=\Theta(1/n)$. Specifically, assume the potential is induced by a univariate function $\phi(z)=\frac{1}{1+\omega}(\psi(z)+\omega z^{2})$. There exists $\omega_{0}>0$ (depending on $\psi$, the training data and the initialization distribution) such that for any $\omega>\omega_{0}$, with high probability over the random initialization, there exist constant $C_{1},C_{2}>0$ such that for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}$ (8)
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$ that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
Theorem 8 relies on the assumption that $\omega$ in $\phi=\frac{1}{1+\omega}(\psi+\omega x^{2})$ is larger than a threshold $\omega_{0}$. In Proposition 27, we give a closed formula for the threshold $\omega_{0}$. Further, numerical experiments in Section 4 indicate the theorem holds even when $\omega$ takes relatively small values, for instance, $\omega=1$.
@@ 3.3 General framework for the proof of the main results
We present a general framework to derive our main results, which builds on the strategy of Jin & Montúfar (2023) with a few key differences. First, we examine the training dynamics of mirror flow using a modified approach compared to gradient flow analysis. Specifically, in the scaled potential case where training may occur outside the kernel regime, we examine how the flatness of the Hessian of the potential affects training. Second, we introduce and analyze the minimal representation cost problem to derive the function description of the implicit bias. This allows us to remove a previous requirement to linearly adjust the data when the network does not include skip connections. The detailed proofs for Theorems 2 and 8 are presented in Appendix A and Appendix B, respectively.
@@
The first step of the proof is to show that, under mirror flow, functions obtained by training the network are close to those obtained by training only the output weights. Let $\tilde{\theta}(t)=\operatorname{vec}(\mathbf{w}(0),\mathbf{b}(0),\tilde{\boldsymbol{a}}(t),d(0))$ denote the parameter under the update rule where $\mathbf{w},\mathbf{b},d$ are kept at their initial values and
@@
Here, we write $\Phi(\tilde{\boldsymbol{a}})$ for the potential function applied to the output weights. For unscaled potentials, $\Phi(\tilde{\boldsymbol{a}})=\sum_{k=1}^{n}\phi(\tilde{a}_{k}-\hat{a}_{k})$; for scaled potentials, $\Phi(\tilde{\boldsymbol{a}})=n^{-2}\sum_{k=1}^{n}\phi(n(\tilde{a}_{k}-\hat{a}_{k}))$. Let $\tilde{\mathbf{f}}(t)=[f(x_{1},\tilde{\theta}(t)),\ldots,f(x_{m},\tilde{\theta}(t))]^{T}\in\mathbb{R}^{m}$. The evolution of $\tilde{\mathbf{f}}(t)$ can be given by
$\frac{\mathrm{d}}{\mathrm{dt}}\tilde{\mathbf{f}}(t)$ $=-n\eta\tilde{H}(t)(\tilde{\mathbf{f}}(t)-\mathbf{y}),\quad\tilde{\mathbf{f}}(0)=\mathbf{f}(0),$ (11)
where $\text{where }\tilde{H}(t)$ $\triangleq n^{-1}J_{\tilde{\boldsymbol{a}}}\tilde{\mathbf{f}}(t)(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\tilde{\boldsymbol{a}}}\tilde{\mathbf{f}}(t)^{T}.$
We aim to show that for any fixed $x\in\mathbb{R}$, $\lim_{n\rightarrow\infty}\sup_{t\geq 0}|f(x,\theta(t))-f(x,\tilde{\theta}(t)|=0$. It suffices to show the kernel matrix $H(t)$ remains sufficiently close to $\tilde{H}(t)$ during training. This approach is well established for gradient flow (Du et al., 2019; Lee et al., 2019) and translates naturally to the case of unscaled potentials. Specifically, we show $\|\theta(t)-\hat{\theta}\|_{\infty}$ and $\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}$ can be bounded by $O(n^{-1/2})$ (Proposition 13 and 14), and then show $H(t)$ stays close to $\tilde{H}(t)$ with the continuity of the Hessian map $\nabla^{2}\Phi(\cdot)$ in Assumption 1 (Proposition 15).
For scaled potentials, we first use a similar strategy to show the training is lazy, i.e., $\|\theta(t)-\hat{\theta}\|_{\infty}$ and $\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}$ are bounded by $O(n^{-1})$ (Proposition 24 and Proposition 25). Then we bound the difference between $\{n\theta(t)\}_{t\geq 0}$ and $\{n\tilde{\theta}(t)\}_{t\geq 0}$, by requiring the Hessian map to be sufficiently flat over a compact region, and then obtain the closeness between $H(t)$ and $\tilde{H}(t)$ (Proposition 26). Further, we show for potential of the form $\phi(x)=\frac{1}{1+\omega}(\psi(x)+\omega x^{2})$, the desired flatness can be achieved by selecting a $\omega$ that is larger than a threshold value $\omega_{0}$ (Proposition 27).
@@
$g(x,\alpha)=\int\alpha(w,b)\sigma(wx-b)\mathrm{d}\mu,\quad\alpha\in\Theta.$ (13)
Here $\Theta\subset C(\operatorname{supp}(\mu))$ is chosen as the parameter space for the model (13). Our goal is to identify a cost functional on the parameter space $\mathcal{C}_{\Phi}\colon\Theta\rightarrow\mathbb{R}$ such that, as the width of the network goes to infinity, the solution to (12) converges to the solution to the following problem: [^3] The convergence is in the following sense: let $\bar{\boldsymbol{a}}$ be the solution to (12) and $\bar{\alpha}_{n}\in\Theta$ be a continuous function satisfying $\bar{\alpha}_{n}(\hat{w}_{k},\hat{b}_{k})=n(\bar{a}_{k}-\hat{a}_{k})$. Let $\bar{\alpha}$ be the solution to (14). Then for any given $x\in\mathbb{R}$, $g_{n}(x,\bar{\alpha}_{n})$ converges in probability to $g(x,\bar{\alpha})$ as $n\rightarrow\infty$.
@@
The last step of our proof is to translate the implicit bias in parameter space, as described by problem (14), to the function space =$\mathcal{F}=\{h(x)=\int\alpha(w,b)\sigma(wx-b)\mathrm{d}\mu\colon\alpha\in\Theta\}$. We do so by “pulling back” the cost functional $\mathcal{C}_{\Phi}$ from the parameter space $\Theta$ to the function space $\mathcal{F}$, via the map $\mathcal{R}_{\Phi}\colon\mathcal{F}\rightarrow\Theta$ defined by
@@
The optimization problem in (17) is referred to as the minimal representation cost problem, as it seeks the parameter $\alpha$ with the minimal cost $\mathcal{C}_{\Phi}(\alpha)$ such that $g(\cdot,\alpha)$ represents a given function $h$. Savarese et al. (2019) utilized this to study the function represented by an infinitely wide ReLU network that interpolates the training data and whose parameters minimize a particular cost $\mathcal{C}(\alpha)=\|\alpha\|_{1}$. In the following result, we show that for general cost functionals, the map $\mathcal{R}_{\Phi}$ allows one to translate problem (14) to a variational problem on the function space. The proof is presented in Appendix A.3.
@@ 4 Experiments
Figure 1: Illustration of Theorem 2. Left: ReLU networks with $4860$ hidden neurons, uniformly initialized input biases and zero initialized output weights and biases, trained with mirror flow on a common data set using unscaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$. Middle: $L_{\infty}$-error between the solution to the variational problem and the networks trained using mirror descent with $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, against the network width. Right: ReLU networks trained with gradient descent on five different data sets, each obtained by translating the same data set along the $y$-axis.
@@
Figure 2: Illustration of Theorem 8. Left: absolute value networks with $4860$ hidden neurons, uniformly initialized input biases and zero initialized output weights and biases, trained with mirror descent on a common data set using scaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$. For comparison we also plot a network with Gaussian initialized input biases, trained with gradient descent. Middle: $L_{\infty}$-error between the solution to the variational problems and networks trained using mirror descent for scaled potentials $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, against the network width. Right: the distribution of the magnitude of the second derivative of the solutions to the variational problems for $\phi_{1}$ (blue), $\phi_{2}$ (orange), and $\phi_{3}$ (green). The inset shows the second derivatives over the input domain.
We illustrate Theorem 8 numerically in Figure 2. The left panel illustrates the effect of the scaled potentials and parameter initialization on the implicit bias of mirror descent. Specifically, different potentials result in functions with different curvature magnitudes, whereas different parameter initializations result in different curvature locations. The middle plot verifies that the solution to the variational problem captures the trained network solutions as the network width increases. The right panel shows the magnitude of the second derivative of the solutions to the variational problems for the potentials in the left panel. We see $\phi_{1}$ allows occurrences of more extreme small and large magnitudes, whereas $\phi_{3}$ allows the second derivative to concentrate more on moderate values.
@@ 5 Conclusion
We focused on univariate, shallow networks with unit-norm initialized input weights. Characterizing the implicit bias for deep nonlinear networks remains an open challenge even for the case of ordinary gradient descent. Extending the results to multivariate networks with general input weights initialization requires stronger theoretical tools (see Appendix E.2 for details). We considered separable potentials which allow for a convenient definition in the infinite width limit and we did not cover potentials defined on a proper subset of the parameter space, such as the negative entropy. Extending our results to more general potentials could be an interesting direction for future work. Other future research directions include deriving closed-form solutions for the variational problems and investigating the rate of convergence as the number of neurons increases.
@@
Code to reproduce our experiments is made available at https://anonymous.4open.science/r/implicit-bias-mirror-descent.

### v2

@@ 3.2 Implicit bias of mirror flow with scaled potentials
Assume the potential function $\Phi\colon\mathbb{R}^{p}\rightarrow\mathbb{R}$ satisfies: (i) $\Phi$ can be written in the form of $\Phi(\theta)=\frac{1}{n^{2}}\sum_{k=1}^{p}\phi\big{(}n(\theta_{k}-\hat{\theta}_{k})\big{)}$, where $\phi$ is a real-valued function on $\mathbb{R}$; (ii) $\phi$ takes the form of $\phi(x)=\frac{1}{1+\omega}\big{(}\psi(x)+\omega x^{2}\big{)}$, where $\omega>0$ and $\psi\in C^{3}(\mathbb{R})$ is a convex function on $\mathbb{R}$.
@@
The following result characterizes the implicit bias of mirror flow with scaled potentials for wide networks with absolute value activations. We consider absolute value activations because, compared to ReLU they yield a smaller function space that makes the problem more tractable.
@@
Consider a two layer network (1) with absolute value activation, $d\geq 1$ input units and $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Additionally, assume $\mathcal{B}$ is compactly supported. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ with $x_{i}\neq x_{j}$ when $i\neq j$ and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with a potential satisfying Assumptions 8 and learning rate $\eta=\Theta(1/n)$. Specifically, assume the potential is induced by a univariate function $\phi(z)=\frac{1}{1+\omega}(\psi(z)+\omega z^{2})$. There exists $\omega_{0}>0$ (depending on $\psi$, the training data and the initialization distribution) such that for any $\omega>\omega_{0}$, there exist constants $C_{1},C_{2}>0$ such that with high probability over the random parameter initialization, for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}$ (8)
Moreover, assuming univariate input data, i.e., $d=1$, and letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$ that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
Theorem 9 relies on the assumption that $\omega$ in $\phi=\frac{1}{1+\omega}(\psi+\omega x^{2})$ is larger than a threshold $\omega_{0}$. In Proposition 30, we give a closed-form formula for this threshold. Further, numerical experiments in Section 4 indicate the theorem holds even when $\omega$ takes relatively small values, for instance, $\omega=1$.
@@ 3.3 General framework for the proof of the main results
We present a general framework to derive our main results, which builds on the strategy of Jin & Montúfar (2023) with a few key differences. First, we examine the training dynamics of mirror flow, for which the analysis is more intricate than that of gradient flow due to the need to account for the evolution of the inverse Hessian of the potential throughout training. In particular, in the scaled potential case, we carefully analyze how the Hessian’s properties, such as the decay rate of its smallest eigenvalue, affect the parameter trajectories. Second, we solve the minimal representation cost problem to derive the function description of the implicit bias. This allows us to remove a previous requirement to linearly adjust the data when the network does not include skip connections. The detailed proofs for Theorems 2 and 9 are presented in Appendix A and Appendix B, respectively.
@@
The first step of the proof is to show that, under mirror flow, functions obtained by training the network are close to those obtained by training only the output weights. Let $\tilde{\theta}(t)=\operatorname{vec}(\boldsymbol{W}(0),\boldsymbol{b}(0),\tilde{\boldsymbol{a}}(t),d(0))$ denote the parameter under the update rule where $\boldsymbol{W},\boldsymbol{b},d$ are kept at their initial values and
@@
Here, we write $\Phi(\tilde{\boldsymbol{a}})$ for the potential function applied to the output weights. For unscaled potentials, $\Phi(\tilde{\boldsymbol{a}})=\sum_{k=1}^{n}\phi(\tilde{a}_{k}-\hat{a}_{k})$; for scaled potentials, $\Phi(\tilde{\boldsymbol{a}})=n^{-2}\sum_{k=1}^{n}\phi(n(\tilde{a}_{k}-\hat{a}_{k}))$. Let $\tilde{\boldsymbol{f}}(t)=[f(x_{1},\tilde{\theta}(t)),\ldots,f(x_{m},\tilde{\theta}(t))]^{T}\in\mathbb{R}^{m}$. The evolution of $\tilde{\boldsymbol{f}}(t)$ can be given by
$\frac{\mathrm{d}}{\mathrm{dt}}\tilde{\boldsymbol{f}}(t)$ $=-n\eta\tilde{H}(t)(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}),\quad\tilde{\boldsymbol{f}}(0)=\boldsymbol{f}(0),$ (11)
where $\text{where }\tilde{H}(t)$ $\triangleq n^{-1}J_{\tilde{\boldsymbol{a}}}\tilde{\boldsymbol{f}}(t)(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\tilde{\boldsymbol{a}}}\tilde{\boldsymbol{f}}(t)^{T}.$
We aim to show that for any fixed $x\in\mathbb{R}^{d}$, $\lim_{n\rightarrow\infty}\sup_{t\geq 0}|f(x,\theta(t))-f(x,\tilde{\theta}(t)|=0$. For unscaled potentials, we adapt an established approach from gradient flow analysis (Du et al., 2019; Lee et al., 2019). Specifically, we show $\|\theta(t)-\hat{\theta}\|_{\infty}$ and $\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}$ are bounded by $O(n^{-1/2})$ (Proposition 14). Then we bound $\|H(t)-\tilde{H}(t)\|_{2}$ with the continuity of the Hessian map $\nabla^{2}\Phi(\cdot)$ in Assumption 1, which gives the desired bound on $|f(x,\theta(t))-f(x,\tilde{\theta}(t)|$ (Proposition 16). For scaled potentials, we use a similar strategy to show $\|\theta(t)-\hat{\theta}\|_{\infty}$ and $\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}$ are bounded by $O(n^{-1})$ (Proposition 27). Then we show $\|n\theta(t)-n\tilde{\theta}(t)\|_{\infty}$ can be bounded when the smallest and largest eigenvalues of $\nabla^{2}\Phi(\theta)$ have mild decay and growth rates, respectively, as $\|\theta\|_{2}$ increases, which allows us to control $|f(x,\theta(t))-f(x,\tilde{\theta}(t))|$ (Proposition 29).
@@
$g(x,\alpha)=\int\alpha(w,b)\sigma(w^{T}x-b)\mathrm{d}\mu,\quad\alpha\in\Theta.$ (13)
Here $\Theta\subset C(\operatorname{supp}(\mu))$ is chosen as the parameter space for the model (13). Our goal is to identify a cost functional on the parameter space $\mathcal{C}_{\Phi}\colon\Theta\rightarrow\mathbb{R}$ such that, as the width of the network goes to infinity, the solution to (12) converges to the solution to the following problem: [^3] The convergence is in the following sense: let $\bar{\boldsymbol{a}}$ be the solution to (12) and $\bar{\alpha}_{n}\in\Theta$ be a continuous function satisfying $\bar{\alpha}_{n}(\hat{w}_{k},\hat{b}_{k})=n(\bar{a}_{k}-\hat{a}_{k})$ for all $k\in[n]$. Let $\bar{\alpha}$ be the solution to (14). Then for any given $x\in\mathbb{R}^{d}$, $g_{n}(x,\bar{\alpha}_{n})$ converges in probability to $g(x,\bar{\alpha})$ as $n\rightarrow\infty$.
@@
The last step of our proof is to translate the implicit bias in parameter space, as described by problem (14), to the function space of univariate networks, =$\mathcal{F}=\{h(x)=\int\alpha(w,b)\sigma(wx-b)\mathrm{d}\mu\colon\alpha\in\Theta\}$. We do so by “pulling back” the cost functional $\mathcal{C}_{\Phi}$ from the parameter space to the function space, via the map $\mathcal{R}_{\Phi}\colon\mathcal{F}\rightarrow\Theta$ defined by
@@
The optimization problem in (17) is known as the minimal representation cost problem, as it seeks the parameter $\alpha$ with the minimal cost $\mathcal{C}_{\Phi}(\alpha)$ such that $g(\cdot,\alpha)$ represents a given function $h$. In the following result, we show that for general cost functionals, the map $\mathcal{R}_{\Phi}$ allows one to translate problem (14) to a variational problem on the function space. The proof is presented in Appendix A.3.
@@ 4 Experiments
Figure 1: Illustration of Theorem 2. Left: ReLU networks with $4860$ hidden neurons, uniformly initialized input biases and zero initialized output weights and biases, trained with mirror flow on a common data set using unscaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$. Middle: $L^{\infty}$-error between the solution to the variational problem and the networks trained using mirror descent with $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, against the network width. Right: ReLU networks trained with gradient descent on five different data sets, each obtained by translating the same data set along the $y$-axis.
@@
Figure 2: Illustration of Theorem 9. Left: absolute value networks with $4860$ hidden neurons, uniformly initialized input biases and zero initialized output weights and biases, trained with mirror descent on a common data set using scaled potentials: $\phi_{1}=x^{2}$, $\phi_{2}=|x|^{3}+x^{2}$, and $\phi_{3}=x^{4}+x^{2}$. For comparison we also plot a network with Gaussian initialized input biases, trained with gradient descent. Middle: $L^{\infty}$-error between the solution to the variational problems and networks trained using mirror descent for scaled potentials $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$, against the network width. Right: the distribution of the magnitude of the second derivative of the solutions to the variational problems for $\phi_{1}$ (blue), $\phi_{2}$ (orange), and $\phi_{3}$ (green). The inset shows the second derivatives over the input domain.
We illustrate Theorem 9 numerically in Figure 2. The left panel illustrates the effect of the scaled potentials and parameter initialization on the implicit bias of mirror descent. When fixing the potential, a network with Gaussian initialization (dashed blue curve) learns a function whose curvature is more concentrated around the origin, compared to a network with uniform initialization (solid blue curve). When fixing the initialization, different potentials result in functions with different curvature magnitudes (solid curves in different colors). The middle plot verifies that the solution to the variational problem captures the trained network solutions as the network width increases. The right panel further illustrates how the potentials in the left panel lead to functions with different curvature magnitudes. We see $\phi_{1}$ allows occurrences of more extreme small and large magnitudes, whereas $\phi_{3}$ allows the second derivative to concentrate more on moderate values.
@@ 5 Conclusion
Our variational characterization of the implicit bias only applies to univariate, shallow networks with unit-norm initialized input weights. Characterizing the implicit bias for deep nonlinear networks remains an open challenge even for ordinary gradient descent. Extending the results to multivariate networks with general input weights initialization requires stronger theoretical tools (see Appendix E.2 for details). We considered separable potentials which allow for a convenient definition in the infinite width limit and we did not cover potentials defined on a proper subset of the parameter space, such as the negative entropy. Extending our results to more general potentials could be an interesting direction for future work. Other future research directions include deriving closed-form solutions for the variational problems and investigating the rate of convergence as the number of neurons increases.
@@
Code to reproduce our experiments is made available at https://github.com/shuangliang15/implicit-bias-mirror-descent.